# Small Bottle, Big Pipe: Quantifying and Addressing the Impact of Data Centers on Public Water Systems

### 03 Mar 2026 v1

@@ 4.2.1 Annual water withdrawal
When benchmarking against total public water supply, we use the baseline values in the most study [49] without adjusting for future system-wide changes. Nationally, under the Baseline scenario, total U.S. data center water withdrawals in 2030 are projected to represent approximately 0.6% to 1.1% of total public water withdrawals in the low- and high-growth cases, respectively. Specifically, without significant improvements in water efficiency and in the high-growth case, total U.S. data center water withdrawals in 2030 could rival the current annual water supply of Los Angeles (from 2019 to 2024), which serves approximately 4 million residents [108].
@@ 5 Recommendations
The lack of technical and financial capacity remains a critical challenge for many U.S. public water systems [28]. Recently, corporate-community partnerships have emerged as a promising approach [3, 91, 19], playing a key role in supporting resilient water infrastructure in host communities. By providing funding or technical assistance for local water system upgrades, data centers can help mitigate the added strain their operations place on public water supplies. Additionally, sourcing non-potable water for industrial use, such as recycled or reclaimed wastewater, can help reduce demand on municipal drinking water supplies for both data center operations and community needs [109]. We therefore recommend that data centers actively participate in these partnerships, aligning corporate responsibility with community well-being and long-term water resilience.
@@ A.6 Peaking Factor
To estimate MDD, a daily peaking factor is required to convert ADD into the water capacity demand measured in terms of the maximum daily demand. In Section 3.1.2, we observe that cooling towers exhibit a relatively lower peaking factor (measured or estimated at approximately 2.0–2.5, which will be higher for planning purposes to account for the worst case [123]), whereas dry cooling with evaporative assistance shows a substantially higher peaking factor due to the concentration of water use on a limited number of days each year. For one leading technology company, the daily peaking factor is estimated at approximately 6.5 (based on the measured monthly peadking factor of 4.3) in West Des Moines, Iowa, planned at 8.0+ in Leesburg, Virginia, and planned at 30+ in Wisconsin. Likewise, a leading technology company’s state-of-the-art data center under construction to host AI and core products has secured an allocation of up to 8 MGD of water capacity, resulting in a peaking factor exceeding 6.3 based on the company’s existing U.S.-wide average water withdrawal intensity (see Appendix D for details).
@@ Appendix C Water Capacity Need for Cooling 100 MW IT Loads
Next, we examine dry cooling with evaporative assistance. While this approach significantly reduces annual WUE compared to conventional cooling towers, peak water demand can still be substantial. As a real example, consider a leading technology company’s data center campus in West Des Moines, Iowa. In 2022, the reported annual WUE was 0.19 L/kWh, [130]. The peak daily WUE (pWUE) can be estimated by multiplying the monthly peaking factor of approximately 3 in 2022 (as shown in Figure 2a) by a minimum adjustment factor of 1.5 based on regulatory guidance [83], yielding a pWUE of 0.855 L/kWh. This is substantially higher than the annual average. Applying the default consumptive ratio of 0.75 [47] results in a water withdrawal intensity of 1.14 L/kWh. Thus, in West Des Moines, Iowa, cooling a 100 MW IT load requires approximately 0.72 MGD, which also needs to be further adjusted upward to approximately 1.0 MGD for planning to maintain operational safety margins in practice.
@@ Appendix E Water-Power Tradeoff for Data Center Cooling
While the high power density of AI server racks often necessitates liquid cooling (e.g., direct-to-chip and immersion cooling) because of its superior heat removal capability [57, 55], they still have approximately 20% or more loads cooled by air that have lower temperature setpoint than liquid-cooled IT loads [15, 139]. This means that evaporative assistance can have significant benefits of peak power reduction for the air-cooled portion of IT loads. For example, to balance the peak power demand and the host community’s limited available water capacity, a large data center uses dry coolers for its liquid-cooled AI servers while evaporative assitance for other servers [138]. Indeed, a leading technology company acknowledges that “water is the most efficient means of cooling in many locations” [17].
@@ Appendix F Additional Recommendations
While water storage tanks can serve as a buffer to smooth hourly water demand peaks, it is challenging to use storage to fully offset multi-day peak water demand during prolonged heatwaves, as this requires costly utility-scale tanks, sanitation, maintenance, and other operational considerations. Nonetheless, data center workloads often exhibit substantial scheduling flexibility [141, 142], which, when combined with multi-region site deployment, may effectively reduce peak water demand. Therefore, in addition to power grid demand response, “water-aware” computing presents a promising approach to lowering peak water demand and strengthening public water infrastructure resilience.

### 18 Mar 2026 v2

@@ 4.2.1 Annual water withdrawal
When benchmarking against total public water supply, we use the baseline values in the most recent study [52] without adjusting for future system-wide changes. Nationally, under the Baseline scenario, total U.S. data center water withdrawals in 2030 are projected to represent approximately 0.6% to 1.1% of total public water withdrawals in the low- and high-growth cases, respectively. Specifically, without significant improvements in water efficiency and in the high-growth case, total U.S. data center water withdrawals in 2030 could rival the current annual water supply of Los Angeles (from 2019 to 2024), which serves approximately 4 million residents [118].
@@ 5 Recommendations
The lack of technical and financial capacity remains a critical challenge for many U.S. public water systems [28]. While some state laws require large water users to make “pro rata” fair contributions to fund water infrastructure expansion [32, 33], additional corporate-community partnerships that go beyond these legal requirements may further strengthen the resilience of water infrastructure in host communities. By providing funding or technical assistance for local water system upgrades, data centers can help mitigate the added strain their operations place on public water supplies. Additionally, sourcing non-potable water for industrial use, such as recycled or reclaimed wastewater, can help reduce demand on municipal drinking water supplies for both data center operations and community needs [119]. We therefore recommend that data centers actively participate in these partnerships, aligning corporate responsibility with community well-being and long-term water resilience.
@@ A.6 Peaking Factor
To estimate MDD, a daily peaking factor is required to convert ADD into the water capacity demand measured in terms of the maximum daily demand. In Section 3.1.2, we observe that cooling towers exhibit a relatively lower peaking factor (measured or estimated at approximately 2.0–2.5, which will be higher for planning purposes to account for the worst case [32]), whereas dry cooling with evaporative assistance shows a substantially higher peaking factor due to the concentration of water use on a limited number of days each year. For one leading technology company, the daily peaking factor is estimated at approximately 6.5 (based on the measured monthly peaking factor of 4.3) in Iowa, planned at 8.0+ in Leesburg, Virginia, and planned at 30+ in Wisconsin. Likewise, a leading technology company’s state-of-the-art data center under construction to host AI and core products has secured an allocation of up to 8 MGD of water capacity, resulting in a peaking factor exceeding 6.3 based on the company’s existing U.S.-wide average water withdrawal intensity (see Appendix D for details).
@@ Appendix C Water Capacity Need for Cooling 100 MW IT Loads
Next, we examine dry cooling with evaporative assistance. While this approach significantly reduces annual WUE compared to conventional cooling towers, peak water demand can still be substantial. As a real example, consider a leading technology company’s data center campus in Iowa. In 2022, the reported annual WUE was 0.19 L/kWh, [139]. The peak daily WUE (pWUE) can be estimated by multiplying the monthly peaking factor of approximately 3 in 2022 (as shown in Figure 2a) by a minimum adjustment factor of 1.5 based on regulatory guidance [92], yielding a pWUE of 0.855 L/kWh. This is substantially higher than the annual average. Applying the default consumptive ratio of 0.75 [50] results in a water withdrawal intensity of 1.14 L/kWh. Thus, for this Iowa data center campus, cooling a 100 MW IT load requires approximately 0.72 MGD, which also needs to be further adjusted upward to approximately 1.0 MGD for planning to maintain operational safety margins in practice.
@@ Appendix E Water-Power Tradeoff for Data Center Cooling
While the high power density of AI server racks often necessitates liquid cooling (e.g., direct-to-chip and immersion cooling) because of its superior heat removal capability [60, 58], they still have approximately 20% or more loads cooled by air that have lower temperature setpoint than liquid-cooled IT loads [15, 148]. This means that evaporative assistance can have significant benefits of peak power reduction for the air-cooled portion of IT loads. For example, to balance the peak power demand and the host community’s limited available water capacity, a large data center uses dry coolers for its liquid-cooled AI servers while evaporative assitance for other servers [147]. Indeed, a leading technology company acknowledges that “water is the most efficient means of cooling in many places” [17].
@@ Appendix F Additional Recommendations
Water storage tanks can serve as a buffer to smooth hourly water demand peaks. However, it is challenging to use storage to fully offset multi-day or even month-long peak water demand during prolonged heatwaves, as this requires costly large-scale tanks, sanitation, maintenance, and other operational considerations. Data center workloads often exhibit substantial scheduling flexibility [150, 151], which, when combined with multi-region site deployment, may effectively reduce peak water demand. Therefore, “water-aware” computing in combination with water tanks presents a promising approach to lowering peak water demand and strengthening public water system resilience.